OpenStudy (anonymous):

solve for p 64p^2-81=0

OpenStudy (anonymous):

\[64p^2-81=(8p+9)(8p-9)=0\] so, \[p=\pm \frac{9}{8}\]

OpenStudy (anonymous):

(8p+9)(8p-9)=0 p=-9/8 p=9/8